OpenStudy (anonymous):

Solve for x: \[\log_{6}(x+4)-\log_{6}(x-4)=\log_{6}3 \]

OpenStudy (anonymous):

@satellite73 @phi @amistre64

OpenStudy (amistre64):

\[log~a-log~b=log~\frac ab\] should be useful

OpenStudy (amistre64):

also:\[if~log_n(p)=log_n(q)~then~p=q\]

OpenStudy (anonymous):

so it looks like \[\log_{6}\frac{ x+4 }{ x-4 } =\log_{6}3 \]

OpenStudy (anonymous):

as the first step..

OpenStudy (phi):

yes, now take the anti-log of both sides then multiply both sides by (x-4)

OpenStudy (phi):

take the anti-log means drop the logs or, mathematically, make each side the exponent of 6 and 6 ^ (log_6(stuff)) = stuff

OpenStudy (anonymous):

i got 8 as my final answer.. is this right?

zepdrix (zepdrix):

yah that sounds right, good job!

OpenStudy (anonymous):

ok cool so you got the same answer then?

zepdrix (zepdrix):

yah

OpenStudy (anonymous):

great! i feel better then haha

OpenStudy (phi):

It is easy to check your answer. Start with the original question \[ \log_{6}\frac{ x+4 }{ x-4 } =\log_{6}3 \] replace x with 8 \[ \log_{6}\frac{ 8+4 }{ 8-4 } =\log_{6}3 \] simplify \[ \log_{6}\frac{ 12 }{ 4 } =\log_{6}3 \] \[ \log_{6}3 =\log_{6}3 \] which looks good...